OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 34 inches. Find the length and width of the rectangle. Find the width of the rectangle. 6 in. 6.5 in. 7.5 in. C

OpenStudy (jack1):

isuru let kesh finish...?

OpenStudy (anonymous):

let l=w+4 therefore,perimeter=2xlxw =2(w+4)w =2(w^2)+8w since we already know that perimeter=34 we have 2w^2 +8w=34 dividing throughout by 2; w^2+4w-17=0 solve the quadratic using the quadratic formula to obtain the value of w

OpenStudy (jack1):

sorry kesh... perimeter not area dude

OpenStudy (anonymous):

were finding the width lol

OpenStudy (***[isuru]***):

hey,  now solve it for x.... I gave kesh a try -->Jack :P